OpenStudy (anonymous):

What is the solution to the equation? 2-z/3=12 A. z = –42 B. z = –34 C. z = –30 D. z = 42

OpenStudy (anonymous):

If your solving for z then you want to get z alone on one side. I would suggest getting rid of the 2 first then multiplying the 3. when you have that done flip the signs so z isnt negative.

OpenStudy (anonymous):

umm i dont get it...

OpenStudy (anonymous):

?

OpenStudy (anonymous):

2-z/3=12 -2 -2 -z/3=10 *3 *3 -z=30 z=30 This is how you would set it up. You would first subtract the two then multiply the three. to get -z=30 then you would flip the signs to get your final answer of z=-30

OpenStudy (anonymous):

What is the solution to the equation? 4 + 3s – 9s = 22 A. s = –3 B. s = –1.5 C. s = 1.5 D. s=4 2/3

OpenStudy (anonymous):

thank ya so much umm more help :D

OpenStudy (anonymous):

Haha! You need to put like terms together on this one so you would have to do the equation 3s-9s first, then subtract the 4 from both sides. then divide whatever is next to s to get your final answer.

OpenStudy (anonymous):

c?